The magnetic field through a single loop of wire 11.5 cm in radius and of 8.91 capital omega resistance, changes with time as shown in Fig. 31-34. The uniform magnetic field is perpendicular to the plane of the loop. Calculate the magnitude of the emf in the loop as a function of time in each of the following time intervals:

The magnetic field through a single loop of wire 1

1.t=0 to t=2

2.t=4 to t=6


Help me please!
Answer #1

a)φ=d(BA)/dt=AdB/dt=πr2*dB/dt=π*0.1152*0.5/2=0.0104V

b)φ=d(BA)/dt=AdB/dt=πr2*dB/dt=π*0.1152*0.5/2=0.0104V